OpenStudy (anonymous):

if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 8 -x - 3y = -12

OpenStudy (anonymous):

-x-3y=-12 -x=-12-3y x=12-3y